You are thinking about investing$4,567in your​ friend's landscaping business. Even though you know the investment is risky and you​ can't be​ sure, you expect your investment to be worth$5,712 next year. You notice that the rate for​ one-year Treasury bills is 1%.However, you feel that other investments of equal risk to your​ friend's landscape business offer an expected return of8%for the year.The present value of the return is? What should you​ do?
Answer #1
Future value = present value*(1+ rate)^time
5712 = Present value*(1+0.01)^1
Present value = 5655.45
Using Calculator: press buttons "2ND"+"FV" then assign
FV =-5712
I/Y =1
N =1
PMT = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(0.01,1,,5712,)
Future value = present value*(1+ rate)^time
5712 = 4567*(1+Interest rate/100)^1
Interest rate % = 25.07
Using Calculator: press buttons "2ND"+"FV" then assign
FV =-5712
PV =-4567
N =1
PMT = 0
CPT I/Y
Using Excel
=RATE(nper,pmt,pv,fv,type,guess)
=RATE(Interest rate/100,,-4567,5712,,)

Choose to invest in the landscaping business as it has highest Return%